The slope of a line is 16 and the point (2, 3) is on the line. Express the line in slope-intercept form.

Answer by DrBeeee(684)

You need
(1) y = mx + b
where the constant m is the slope and the constant b is the y-intercept.
Your problem statment provides
(2) m = 16
To get b, substitute the point (2,3) and m into (1) to get
(2) 3 = 16*2 + b
Solve (2) for b
(3) b = -29
Your answer is
(4) y = 16x - 29
To check let (x,y) = (2,3)
Is (3 = 16*2 -29)?
Is (3 = 32 - 29)?
Is (3 = 3)? Yes
There is another technique to obtain (4) from m and the point (2,3).
Use
(5) (y - 3) = 16*(x - 2)
Simplifying (5) yields
(6) y = 16x -29, which is the same as (4)